OpenStudy (anonymous):

PLEASE HELP Use basic identities to simplify the expression 1/ (cot^2 θ )+ sec θ cos θ

OpenStudy (anonymous):

sec^2θ

OpenStudy (anonymous):

i believe...

OpenStudy (zehanz):

Well, since I'm from Europe, I'll convert everything to tan, sin and cos functions ;) \[\frac{ 1 }{ \cot^2 \theta }+\sec \theta \cos \theta =\tan^2 \theta +\frac{ 1 }{ \cos \theta }\cos \theta=\tan^2 \theta + 1\]Looks good enough for me!

OpenStudy (anonymous):

1/cot = tan sec/cos = 1 tan^2 + 1 = sec^2 so your answer i believe is

OpenStudy (anonymous):

Thank you! can you help me with another question real quick? :)

OpenStudy (zehanz):

Just quickly ask it!

OpenStudy (anonymous):

Simplify the expression. (cos^2 x+ sin^2 x)/(cot^2x -csc^2 x)

OpenStudy (zehanz):

You MUST know a real simple answer to cos²x+sin²x!

OpenStudy (zehanz):

While you are thinking about that one, I'll give cot^2x -csc^2 x a try...

OpenStudy (kainui):

The entire thing can be solved by knowing: sin^2+cos^2=1 divide the whole thing by cos^2 cot^2+1=csc^2

OpenStudy (zehanz):

\[\cot^2x -\csc^2 x=\frac{ \cos^2 x }{ \sin^2 x }-\frac{ 1 }{ \sin^2 x }=\frac{ \cos^2x - 1 }{ \sin^2 x }\] is what I have so far... Now that looks promising! I really need to now the answer of cos²x+sin²x now!

OpenStudy (anonymous):

cos²x+sin²x equals 1

OpenStudy (kainui):

Now divide both sides of the equation by sin^2(x). You'll see you can plug in entire equations for the top and bottom.

OpenStudy (zehanz):

OK, thanks @Kainui, I'm sure @haleyw knew that! So: we've got \[\frac{ 1 }{ \frac{ \cos^2 x -1 }{ \sin^2 x } }\]

OpenStudy (kainui):

No, you don't know it either, that's why I'm telling you.

OpenStudy (kainui):

(cos^2 x+ sin^2 x)/(cot^2x -csc^2 x) cos^2 x+ sin^2 x=1 divide this by sin^2(x) cot^2x+1=csc^2x subtract 1 and csc^2(x) from either side cot^2x-csc^2x=-1 Plug them both into the original equation (1)/(-1)

OpenStudy (zehanz):

Just turn upside down:\[\frac{ \sin^2 x }{ \cos^2 -1 }=-\frac{ \sin^2 x }{ 1- \cos^2 x }=-\frac{ \sin^2 x }{ \sin^2 x }=-1\]

OpenStudy (anonymous):

Thank you both so much! I really appreciate it!

OpenStudy (zehanz):

In this last step, I used if cos²x + sin²x=1, then sin²x = 1 - cos²x. I only had cos²x - 1, so I reversed them and put a minus sign in front of the fraction.

OpenStudy (zehanz):

It seems to be customary to express your appreciation by pressing the "Best Response" buttons...;)

OpenStudy (zehanz):

Glad to be of help!

OpenStudy (anonymous):

:)